A force of 2N stretch an elastic material by 30mm what additional force will stretch the material 35mm assume that the elastic limit is not exceeded
1
Expert's answer
2020-06-02T17:51:54-0400

The Hooke's law says

"F=kx"

Hence, the stiffness


"k=\frac{F}{x}=\frac{2\:\rm N}{0.03\:\rm m}=66.7\:\rm N/m"

The additional force

"\Delta F=k\Delta x\\
=66.7\:\rm N/m\times 0.005\:\rm m=0.33\:\rm N"